OpenStudy (dls):

A farmer has a piece of land.He wishes to divide equally in his two sons to maintain peace and harmony in family.If his land is denoted by the area y^2=4x and x=4 and to divide the area equally,he draws a line x=a.What is the value of a?

OpenStudy (dls):

Attempt: \[\LARGE \int\limits_0^2 2\sqrt{x}dx=\int\limits_4^a 2 \sqrt{x}dx\] \[\LARGE a^{\frac{3}{2}}=16\] \[\LARGE a=16^{\frac{2}{3}}\] but the answer is slightly different,dunno why..suggestions?

OpenStudy (dls):

@ganeshie8 !

ganeshie8 (ganeshie8):

im getting a = 16^(1/3)

OpenStudy (dls):

how power 1/3?

ganeshie8 (ganeshie8):

i took first integral 0->a second integral a->4

OpenStudy (dls):

:O

OpenStudy (dls):

 sorry a typo there in my eqn its from 0 to 4 in first integral,why a ?